# Bayesian Model Parameter Learning in Linear Inverse Problems: Application in EEG Focal Source Imaging

### v1

@@ 4.2 Skull Conductivity Estimation
$\Gamma_{\varepsilon_{l}}=\gamma\left(\mathrm{var}(\sigma)J_{1}^{l}{J_{1}^{l}}^{\mathrm{T}}+\sum_{k=2}(\mathrm{cov}[z^{k}])+\mathbb{E}[z^{k}]\mathbb{E}[z^{k}]^{\mathrm{T}})J^{l}_{k}{J^{l}_{k}}^{\mathrm{T}}\right),$ (25)
where $z_{k}=(\sigma-\sigma_{*})^{k}$, $J_{k}^{l}=\frac{\partial^{k}A^{l}(\sigma_{*})}{\partial\sigma^{k}}\in\mathbb{R}^{m}$ is the $k$th derivative of the leadfield $A^{l}(\sigma)$ [^2] In practice, we noticed that the derivatives of order $k\geq 2$ were negligible. and $\gamma$ is the dipole variance. Here, without a loss of generality, we have considered that the standard conductivity is equal to the mean conductivity of the training data $\pi(\sigma)$, i.e. $\sigma_{0}\approx\sigma_{*}$, to ease our analysis.
Based on the previous expression for the error covariance, we see that $x_{i}$ has a scaling effect on the error covariance, and the basis $W_{i}$ from the eigenvalue decomposition of the approximation error covariance matrix are primarily related to the model parameter $\sigma$. So, we can expect that the error coefficients $\alpha$ will be proportional to the amplitude of $x_{i}$. Therefore, we can suspect that the methods that employ only the estimated error $\hat{\alpha}_{l}$ to infer $\sigma$, as the conditional Gaussian approximation described below, will not perform as well as the methods that include both, $\hat{\alpha}_{l}$ and the estimated source $\hat{x}_{l}$ as inputs.
4.2.1 Conditional Gaussian Approximation
@@
where $\sigma_{*}$ is the mean value of the postulated skull conductivity distribution, $\Gamma_{\sigma\alpha_{l}}$ is the cross-covariance between $\sigma$ and $\alpha_{l}$ (estimated from the samples of $\alpha_{l}^{(s)}$ and $\sigma^{(s)}$, as will be described in Section 5.4), and $\Gamma_{\alpha_{l}}=\mathrm{diag}\{\lambda_{1l},\ldots,\lambda_{pl}\}$ where $\lambda_{1p}$ are the first $p$ eigenvalues of covariance $\Gamma_{\varepsilon_{l}}$.
@@
4.2.2 Conditional Gaussian with Iterations
In some cases, it is possible to improve the CG-based model parameter estimates by iteratively updating the linearization point. Algorithm 1 presents an iterative approach that uses the CG result as an initialization and also considers the estimated value of the source.
@@ 4.2.3 Gaussian Process
5 Materials and Methods
5.1 Meshes
For our study, we built the two used meshes (a fine and a coarse one) with the help of the MRI data of the so-called ernie subject and SimNIBS 4 software [^3] https://simnibs.github.io/simnibs/build/html/index.html [53]. Four different tissue compartments (scalp, skull, cerebrospinal fluid, and brain) were considered, and 76 electrodes were placed around the head according to the international 10-10 system (see, Fig.1). The fine mesh consisted of 2,103,623 elements joined in 377,150 nodes, and the coarse mesh of 743,575 elements joined in 136,868 nodes.
5.2 Leadfield Models
@@
First, we created $K=200$ leadfield matrices with skull conductivity samples $\sigma^{(k)}$ drawn from a bounded Gaussian distribution $\pi(\sigma)$ with mean $\sigma_{*}=0.0103$ S/m and standard deviation 0.0035 S/m. The skull conductivity distribution ranged from 0.0041 S/m to 0.033 S/m according to the values reported in [55, 56, 57]. The rest of the tissue electric conductivity values were 0.43 S/m for the scalp, 1.79 S/m for cerebrospinal fluid, and 0.33 S/m for the brain (gray matter and white matter) [58]. We refer to these leadfield matrices as sample models, $A(\sigma)$. Out of these, 150 sample models were used to estimated the skull conductivity related approximation error statistics (in Section 5.4) and the remaining 50 sample models were used for the training of the GP (in Section 5.4.2).
For the final testing, we created two accurate leadfield matrices $A(\sigma)$ with the help of the fine mesh and skull conductivity values 0.0061 S/m and 0.0139 S/m, that were not included in the sample set. In addition, we created a standard head model, $A_{0}$, with the help of the coarse mesh and skull conductivity $\sigma_{0}=0.0103$ S/m.
5.3 Source Spaces
The dipole source space was restricted to an approximately 30 mm thick cross sectional area close to the sensory area (as shown in the left image of Fig. 1). This area was selected because previous studies had found it suitable for model calibration purposes using EEG recordings [59, 60]. The source space of the standard model (in the coarse mesh) is marked with blue circles in the right image of Fig. 1, and it uniformly covers the gray matter of the brain. The red dots in the right image of Fig. 1 indicate the source locations (in the fine mesh) that were used in the estimation of the approximation error statistics, as described in Section 5.4. The number of points (source locations) in both source spaces was 560. For the final testing, to avoid over-fitting, we produced observations from simulated dipole sources placed in different locations (in the fine mesh, marked with yellow in the right image of Fig. 1) than the sources that were used to produce the statistics.
@@ 5.3 Source Spaces
5.4 Approximation Error Statistics
The samples for the estimation of the approximation error statistics at location $i$ were created by first choosing randomly one of the sample models and then evaluating both the sample model and the standard model with the same single dipole source $x_{i}^{(j)}$ at location $i$ as
@@ 5.4 Approximation Error Statistics
A set of $J=100$ single radial dipole samples $x_{i}^{(j)}$ with amplitudes drawn from a Rayleigh distribution $\pi(|x_{i}|)|=\text{Rayleigh}(\sqrt{2}\gamma)$ with parameter $\gamma=1.85$ and mode 2.55, and 150 different sample leadfield models were used to estimate the approximation error statistics at each location $i$. The superscripts are as follows: $s=j+J(k-1)$ where $j=1,\ldots,J$, and $k=1,\ldots K$, and $s=1,\ldots,S$. The mean $\varepsilon_{i*}$ and covariance matrix $\Gamma_{\varepsilon_{i}}$ of the corresponding approximation error at location $i$ were
@@
Samples $\alpha_{i}^{(s)}$ and $\varepsilon_{i}^{\prime\prime(s)}$ were evaluated according to $\alpha_{i}^{(s)}=W_{i}^{\mathrm{T}}(\varepsilon_{i}^{(s)}-\varepsilon_{i*})$ and $\varepsilon_{i}^{\prime\prime(s)}=Q_{i}Q_{i}^{\mathrm{T}}(\varepsilon_{i}^{(s)}-\varepsilon_{i*})$, where $W_{i}$ and $Q_{i}$ were matrices with columns $1,\dots,p$ and $p+1,\dots,m$ of the eigenvectors of $\Gamma_{\varepsilon_{i}}$, respectively. Furthermore, by using these samples we estimated numerically $\Gamma_{\alpha_{i}}$ and the cross-covariance $\Gamma_{\sigma\alpha_{i}}$.
@@ 5.4.1 Selection of $p$ Eigenvectors
The number of eigenvectors $p$ required to describe the primary error $\varepsilon^{\prime}$ depends on the discrepancies that produce the approximation errors. In general, it is preferable if only few eigenvectors can be used. In the current setup, when the source location was fixed, the source amplitude had only a scaling effect on the covariance matrix $\Gamma_{\varepsilon_{i}}$. Furthermore, since we were looking for a single model parameter value (skull conductivity $\sigma>0$), we could set $p=1$ and use only the eigenvector corresponding to the largest eigenvalue. Thus, we had a single error coefficient $\alpha_{i}\in\mathbb{R}$.
@@ 5.4.2 Inference of Skull Conductivity from Training Data
Figure 2: Left: For a given source location $i$, we visualize here the sample pairs $\{\alpha_{i}^{(s)},\sigma^{(s)})\}$ for different source amplitudes. The red line shows the conditional conductivity $\sigma_{*|\alpha_{i}}=\sigma_{*}+\Gamma_{\sigma\alpha_{i}}\Gamma_{\alpha_{i}}^{-1}\alpha_{i}^{(s)}$ (26). The green cross shows the sample mean $\sigma_{*}$ and the mean of the error coefficient. Right: Here, we show how the error coefficient $\alpha_{i}$ is related to the skull conductivity value $\sigma$ for a fixed source amplitude. In this case, we can observe monotonic relationships between conductivity and $\alpha_{i}$.
@@
To allow more complex relationships (instead of just linear) and to take into account the effect of the source amplitude, we employd a Gaussian process to infer the skull conductivity. For this purpose, we needed to define the mean and covariance functions, specified by hyperparameters. In the current problem, the statistics of the underlying (unknown) function $\sigma=f(\alpha_{i},x_{i})$, $f\sim\mathcal{N}(m(\alpha_{i},x_{i}),K_{\alpha_{i},x_{i}})$, could be approximated based on the expected relationships between the parameters and simulated training data. In particular, the mean function $m(\alpha_{i},x_{i})$ could be modeled as a polynomial of the following form
$m(\alpha_{i},x_{i})=\sum_{0\leq t\leq T}c_{t}\,\left(\frac{\alpha_{i}+k}{|x_{i}|}\right)^{t},$ (30)
where $k=w_{i1}^{\mathrm{T}}\varepsilon_{i*}$. The covariance kernel was modeled using the following exponential function
$K_{\alpha_{i},x_{i}}=s_{f}^{2}\exp{\left(-\frac{1}{L^{2}}\begin{bmatrix}\frac{\alpha_{i}}{|x_{i}|}-\frac{\alpha_{i}^{\prime}}{|x_{i}^{\prime}|}\end{bmatrix}\begin{bmatrix}\frac{\alpha_{i}}{|x_{i}|}-\frac{\alpha_{i}^{\prime}}{|x_{i}^{\prime}|}\end{bmatrix}^{T}\right),}$ (31)
@@
Figure 3: The blue dots are the target values of the conductivities, the red line is the mean of the GP prediction, the gray color shows the 95% confidence intervals, and the red crosses depict the training data.
In Fig. 3, the blue dots show the simulated (target) values that were not used in the estimation of the $c_{t}$ coefficient, the red line shows the mean prediction of the Gaussian process regression given by (30), the shaded gray areas represent the (pointwise) two standard deviation limits corresponding to each input, and the red crosses show the observations / training data. Here, we can see that the order $T=2$ is able to describe the relationships between the three parameters. Higher order $T$ did not improve the fit.
In practice to apply the GP, we divided the artificially generated training data set $\{x_{i}^{(s)},\alpha_{i}^{(s)},\sigma^{(s)}\}_{s=1}^{N_{s}}$ into two subsets. The first subset was used to estimate the coefficients $c_{t}$ through least squares, and the second subset was used later as input-output training data when solving the problem (27). It is worth noting that the triplets were from additional samples that were not used in the estimation of the approximation error statistics $\Gamma_{\varepsilon_{i}}$, $\varepsilon_{i*}$ nor $W_{i}$ in Section 5.4.
5.5 Testing Data and Different Estimates
@@
5.5.1 Dipole Source Estimates
Since we studied only single dipole sources, we used the single dipole scan algorithm [51] to estimate the sources.
@@ 5.5.1 Dipole Source Estimates
    $(\hat{x}_{l},\hat{\alpha}_{l})\leftarrow\min_{i=1:n}\{\|L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}(v-A_{0}^{i}\hat{x}_{i}-\varepsilon_{i*}-w_{i1}{\alpha}_{i}-e_{*})\|_{2}^{2}+\|L_{\alpha_{i}}{\alpha}_{i}\|_{2}^{2}\},$ (34)

### v2

@@ 4.2 Skull Conductivity Estimation
$\Gamma_{\varepsilon_{l}}=\gamma\left(\mathrm{var}(\sigma)J_{1}^{l}{J_{1}^{l}}^{\mathrm{T}}+\sum_{k,j\geq 2}\mathrm{cov}[z_{k},z_{j}]J^{l}_{k}{J^{l}_{j}}^{\mathrm{T}}\right),$ (25)
where $z_{k}=(\sigma-\sigma_{*})^{k}$, $J_{k}^{l}=\frac{1}{k!}\frac{\partial^{k}A^{l}(\sigma_{*})}{\partial\sigma^{k}}\in\mathbb{R}^{m}$ is the $k$th derivative of the leadfield $A^{l}(\sigma)$ [^3] In practice, we noticed that the derivatives of order $k,j\geq 2$ were negligible. and $\gamma$ is the dipole variance. Here, without a loss of generality, we have considered that the standard conductivity is equal to the mean conductivity of the training data $\pi(\sigma)$, i.e. $\sigma_{0}\approx\sigma_{*}$, to ease our analysis.
From the previous expression for the error covariance, we observe that the source $x_{l}$ introduces a scaling effect on $\Gamma_{\varepsilon_{l}}$. Because of that and since the error covariance is $\Gamma_{\varepsilon_{l}}\approx W^{p}_{l}\Gamma_{\alpha_{l}}(W^{p}_{l})^{\mathrm{T}}$, the bases $W_{l}^{p}$ are shaped primarily by the variability in the model parameter $\sigma$. However, due to the scaling effect of the source $x_{l}$ in $\Gamma_{\varepsilon_{l}}$, one can expect that the error coefficients $\alpha_{l}$ will be proportional to the amplitude of $x_{l}$. Therefore, we can suspect that the methods that employ only the estimated error $\hat{\alpha}_{l}$ to infer $\sigma$, as the conditional Gaussian approximation described below, will not perform as well as the methods that include both, $\hat{\alpha}_{l}$ and the estimated source $\hat{x}_{l}$ as inputs.
4.2.1 Conditional Gaussian (CG) Approximation
@@
where $\sigma_{*}$ is the mean value of the postulated skull conductivity distribution, $\Gamma_{\sigma\alpha_{l}}$ is the cross-covariance between $\sigma$ and $\alpha_{l}$ (estimated from the samples of $\alpha_{l}^{(s)}$ and $\sigma^{(s)}$, as will be described in Section 5.1.4), and $\Gamma_{\alpha_{l}}=\mathrm{diag}\{\lambda_{1l},\ldots,\lambda_{pl}\}$ where $\lambda_{1p}$ are the first $p$ eigenvalues of covariance $\Gamma_{\varepsilon_{l}}$.
@@
4.2.2 Conditional Gaussian with Iterations (CG+Iter.)
In this case, it is possible to improve the CG-based model parameter estimates by iteratively updating the linearization point. Algorithm 1 presents an iterative approach that uses the CG result as an initialization and also considers the estimated value of the source.
@@ 4.2.3 Gaussian Process (GP)
5 Implementation, Results and Discussion
In this Section, we use simulated EEG recordings generated by single-dipole sources within the somatosensory cortex of the brain to explore the capability of the proposed approaches to recover single sources and infer skull conductivity values. In Section 5.1, we present technical details regarding the meshes used in forward simulations and inversion, dipole source spaces, training data, approximation error sampling, and parameter selection procedures in skull conductivity inference. In addition, we describe the used testing data and the different estimates that were computed. Subsequently, in Subsection 5.2, we describe our simulated experiments and present the results under different scenarios to demonstrate the proposed approaches. Finally, we discuss further extensions and applications in Subsection 5.3.
5.1 Implementation Details
5.1.1 Meshes
For our study, we built the two used 3D meshes (a fine and a coarse one) with the help of the MRI data of the so-called ernie subject and SimNIBS 4 software [^4] https://simnibs.github.io/simnibs/build/html/index.html [63]. Four different tissue compartments (scalp, skull, cerebrospinal fluid, and brain) were considered, and 76 electrodes were placed around the head according to the international 10-10 system (see, Fig.1). The fine mesh consisted of 2,103,623 tetrahedral elements joined in 377,150 nodes, and the coarse mesh of 743,575 tetrahedral elements joined in 136,868 nodes.
5.1.2 Leadfield Models
@@
First, we created $K=200$ leadfield matrices with skull conductivity samples $\sigma^{(k)}$ drawn from a bounded Gaussian distribution $\pi(\sigma)$ with mean $\sigma_{*}=0.0103$ S/m and standard deviation 0.0035 S/m. The skull conductivity distribution ranged from 0.0041 S/m to 0.033 S/m according to the values reported in [65, 66, 37]. The rest of the tissue electric conductivity values were 0.43 S/m for the scalp, 1.79 S/m for cerebrospinal fluid, and 0.33 S/m for the brain (gray matter and white matter) [67]. We refer to these leadfield matrices as sample models, $A(\sigma)$. Out of these, 150 sample models were used to estimate the skull conductivity related approximation error statistics (in Section 5.1.4) and the remaining 50 sample models were used for the training of the GP (in Section 5.1.6).
For the final testing, we created two accurate leadfield matrices $A(\sigma)$ with the help of the fine mesh and skull conductivity values 0.0061 S/m and 0.0139 S/m, that were not included in the sample set. In addition, we created a standard model, $A_{0}$, with the help of the coarse mesh and skull conductivity $\sigma_{0}=0.0103$ S/m.
5.1.3 Source Spaces
The dipole source space was restricted to an approximately 30 mm thick cross sectional area close to the somatosensory area (as shown in the left image of Fig. 1). This area was selected because previous studies had found it suitable for model calibration purposes [33, 34, 35, 37, 38]. In the right image of Fig. 1, we show a magnification of the region of interest, highlighting the source spaces corresponding to the models considered. In particular, the source space of the standard model (in the coarse mesh) is marked with blue circles in the right image of Fig. 1, and it uniformly covers the gray matter of the brain with 2.5 mm resolution (distance between neighboring sources). The red dots in the right image of Fig. 1 indicate the source locations (in the fine mesh) that were used in the estimation of the approximation error statistics, as described in Section 5.1.4. The number of points (source locations) in both source spaces was 560. For the final testing, to avoid over-fitting, we produced observations from simulated dipole sources placed in different locations (in the fine mesh, marked with yellow in the right image of Fig. 1) than the sources that were used to produce the statistics.
@@ 5.1.3 Source Spaces
5.1.4 Approximation Error Statistics and Training Data
The approximation error statistics were estimated from samples. These samples for each location index $i$ were created by evaluating one of the (randomly chosen) sample models $A(\sigma^{(k)})$ and the standard model $A_{0}$ with a single dipole source $x_{i}^{(j)}$ as
@@ 5.1.4 Approximation Error Statistics and Training Data
A set of $J=100$ single radial (with respect to the center of the brain) dipole samples $x_{i}^{(j)}$ with amplitudes drawn from a Rayleigh distribution $\pi(|x_{i}|)|=\text{Rayleigh}(\sqrt{2}\gamma)$ with parameter $\gamma=1.85$ and mode 2.55, and 150 different sample leadfield models were used to estimate the approximation error statistics at each location $i$. The superscripts are as follows: $s=j+J(k-1)$ where $j=1,\ldots,J$, and $k=1,\ldots,K$, and $s=1,\ldots,S$. The mean $\varepsilon_{i*}$ and covariance matrix $\Gamma_{\varepsilon_{i}}$ of the corresponding approximation error at location $i$ were
@@
Based on these, we evaluated samples for $\alpha_{i}^{(s)}$ and $\varepsilon_{i}^{\prime\prime(s)}$ as $\alpha_{i}^{(s)}=W_{i}^{\mathrm{T}}(\varepsilon_{i}^{(s)}-\varepsilon_{i*})$ and $\varepsilon_{i}^{\prime\prime(s)}=Q_{i}Q_{i}^{\mathrm{T}}(\varepsilon_{i}^{(s)}-\varepsilon_{i*})$, where $W_{i}$ and $Q_{i}$ contained the eigenvectors $1,\dots,p$ and $p+1,\dots,m$ of $\Gamma_{\varepsilon_{i}}$ respectively (please see Appendix A1 for details). Furthermore, by using these samples we estimated numerically $\Gamma_{\alpha_{i}}$ and the cross-covariance $\Gamma_{\sigma\alpha_{i}}$, i.e. $\Gamma_{\alpha_{i}}=\frac{1}{S-1}\sum_{s=1}^{S}\alpha_{i}^{(s)}(\alpha_{i}^{(s)})^{\mathrm{T}}$, $\Gamma_{\sigma\alpha_{i}}=\frac{1}{S-1}\sum_{s=1}^{S}(\sigma_{i}^{(s)}-\sigma_{i*})(\alpha_{i}^{(s)})^{\mathrm{T}}$ and $\sigma_{i*}=\frac{1}{S}\sum_{s=1}^{S}\sigma_{i}^{(s)}$.
@@ 5.1.5 Selection of $p$ Eigenvectors
The number of eigenvectors $p$ required to describe the primary error $\varepsilon^{\prime}$ depends on the discrepancies that produce the approximation errors. In general, it is preferable if only few eigenvectors can be used. In the current setup, the source amplitude had only a scaling effect on the covariance matrix $\Gamma_{\varepsilon_{i}}$. Furthermore, we noticed that the numerically estimated covariance is well approximated by the first term of equation (25). Then, since we were looking for a single model parameter value (skull conductivity $\sigma>0$), we set $p=1$ and use only the eigenvector corresponding to the largest eigenvalue. Thus, we had a single error coefficient $\alpha_{i}\in\mathbb{R}$ to estimate.
@@ 5.1.6 Inference of Skull Conductivity from Training Data
Figure 2: In these figures, x-axis shows the sampled $\alpha_{i}^{(s)}$ and y-axis the corresponding $\sigma^{(s)}$. Left: For a given source location $i$, we visualize the sample pairs $\{\alpha_{i}^{(s)},\sigma^{(s)})\}$ for all the different source amplitudes. The red line shows the CG skull conductivity predictions $\sigma_{*|\alpha_{i}}=\sigma_{*}+\Gamma_{\sigma\alpha_{i}}\Gamma_{\alpha_{i}}^{-1}\alpha_{i}^{(s)}$ (26). The green cross shows the sample mean $\sigma_{*}$ and the mean of the error coefficient which is zero. Right: Here, we show how the error coefficient $\alpha_{i}^{(s)}$ is related to the skull conductivity value $\sigma^{(s)}$ given a fixed source amplitude $|x_{i}|$. In this case, we can observe monotonic relationships between conductivity $\sigma$ and $\alpha_{i}$.
@@
To allow more complex relationships (instead of just linear) and to take into account the effect of the source amplitude, we employed a Gaussian process to infer the skull conductivity. For this purpose, we needed to define the mean and covariance functions, specified by hyperparameters. In the current problem, the statistics of the underlying (unknown) function $\sigma=f(\alpha_{i},x_{i})$, $f\sim\mathcal{N}(m(\alpha_{i},x_{i}),K_{\alpha_{i},x_{i}})$, could be approximated based on the expected relationships between the parameters and simulated training data. In particular, the mean function $m(\alpha_{i},x_{i})$ could be modeled as a polynomial of the following form
$m(\alpha_{i},x_{i})=\sum_{0\leq{h}\leq{H}}c_{t}\,\left(\frac{\alpha_{i}+k}{|x_{i}|}\right)^{h},$ (30)
where $k=w_{i1}^{\mathrm{T}}\varepsilon_{i*}$. The elements of the covariance kernel were modeled using the following exponential function
$k_{\alpha_{i},x_{i}}=s_{f}^{2}\exp{\left(-\frac{1}{L^{2}}\begin{bmatrix}\frac{\alpha_{i}}{|x_{i}|}-\frac{\alpha_{i}^{\prime}}{|x_{i}^{\prime}|}\end{bmatrix}^{2}\right),}$ (31)
@@
Figure 3: The blue dots are the target values of the conductivities, the red line is the mean of the GP prediction, the gray color shows the two standard deviation intervals, and the red crosses depict the training data.
In Fig. 3, the blue dots show the simulated (target) values that were not used in the estimation of the $c_{t}$ coefficient, the red line shows the mean prediction of the Gaussian process regression given by (30), the shaded gray areas represent the (pointwise) two standard deviation limits corresponding to each input, and the red crosses show the observations / training data. Here, we can see that the order $H=2$ is able to describe the relationships between the three parameters. Higher order $H$ did not improve the fit.
In practice to apply the GP, we divided the artificially generated training data set $\{x_{i}^{(s)},\alpha_{i}^{(s)},\sigma^{(s)}\}_{s=1}^{N_{s}}$ into two subsets. The first subset was used to estimate the coefficients $c_{t}$ through least squares, and the second subset was used later as input-output training data when solving the problem (27). It is worth noting that the triplets were from additional samples that were not used in the estimation of the approximation error statistics $\Gamma_{\varepsilon_{i}}$, $\varepsilon_{i*}$ nor $W_{i}$ in Section 5.1.4.
5.1.7 Testing Data and Different Estimates
@@
Additive Measurement Noise:
Random white noise with zero mean, $e_{*}=0$, was added to the observations. In particular, the measurement noise was $e=\mathrm{s_{c}}\;\bar{e}\in\mathbb{R}^{m}$ where $\bar{e}\in\mathbb{R}^{m}$ was sampled from a Gaussian distribution with zero mean and variance one. The scaling parameter $\mathrm{s_{c}}\in\mathbb{R}^{+}$ was estimated with the help of the chosen signal-to-noise ratio (SNR), defined as $\mathrm{SNR}=10\log_{10}\frac{\|v\|_{2}^{2}}{\|\mathrm{s_{c}}\;\bar{e}\|_{2}^{2}}$ where .$\|.\|_{2}^{2}$ is the $\ell_{2}$- norm.
In the numerical experiments, two levels of SNR were used, 40 dB and 30 dB. In the following, the measurement noise covariance was $\Gamma_{e}=\mathrm{s_{c}}^{2}I_{m}$.
Single Dipole Source Estimates:
Since we studied only single dipole sources, we used the single dipole scanning algorithm [58, 59, 60, 61] to estimate the sources.
@@
    $(\hat{x}_{l},\hat{\alpha}_{l})\leftarrow\min_{i=1:n}\{\|L_{\varepsilon^{{}^{\prime\prime}}_{i}+e}(v-A_{0}^{i}\hat{x}_{i}-\varepsilon_{i*}-w_{i1}{\alpha}_{i})\|_{2}^{2}+\|L_{\alpha_{i}}{\alpha}_{i}\|_{2}^{2}\},$ (34)